Floor plans for a building have the four corners to a room located at points (2,3), (11,6), (-3,18) and (8,21). Determine whether the side through the
points (2,3) and (11,6) is perpendicular to the side through the points (2,3) and (-3,18).

Answer by checkley71(8403)

To solve this problem we need to find the slopes of each of these 2 lines:
(6-3)/(11-2)=3/9=1/3
(18-3)/(-3-2)=15/-5=-3
Because these 2 slopes are negative recipricals of each other then they are perpendicular lines.